# Optimal transport for types and convex analysis for definable predicates in tracial $\mathrm{W}^{*}$-algebras

### 02 Jan 2024 v2

@@ 2.4. Theories, models, and axioms
The class $\mathcal{C}$ is said to be axiomatizable if every $\mathcal{L}$-structure that models $\operatorname{Th}(\mathcal{C})$ is actually in $\mathcal{C}$. Farah, Hart, and Sherman [22, §3.2] showed that the class of $\mathcal{L}_{\operatorname{tr}}$-structures that represent actual tracial $\mathrm{W}^{*}$-algebras is axiomatizable. The theory of tracial $\mathrm{W}^{*}$-algebras will be denoted $\mathrm{T}_{\operatorname{tr}}$. Similarly, the class of $\mathrm{II}_{1}$-factors (infinite-dimensional tracial von Neumann algebras with trivial center) is axiomatizable and we denote its theory by $\mathrm{T}_{\mathrm{II}_{1}}$ [22].
@@ 2.5. Types, definable predicates, and definable functions
The space of types $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is equipped with a weak-*** topology as follows. If $\mathbf{S}$ is an $I$-tuple of $\mathcal{L}$-sorts, the set $\mathcal{F}_{\mathbf{S}}$ of formulas defines a real vector space. For each $\mathcal{L}$-structure $\mathcal{M}$ and $\mathbf{X}\in\prod_{i\in I}S_{j}^{\mathcal{M}}$, the type $\operatorname{tp}^{\mathcal{M}}(\mathbf{X})$ is a linear map $\mathcal{F}_{\mathbf{S}}\to\mathbb{R}$. Thus, for each $\mathcal{L}$-theory $\mathrm{T}$ and $\mathbf{D}\in\prod_{i\in I}\mathcal{D}_{S_{i}}$, the space $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is a subset of the real dual $\mathcal{F}_{\mathbf{S}}^{\dagger}$. We equip $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ with the weak-$\star$ topology (also known as the logic topology), which makes it into a compact Hausdorff space [6, Proposition 8.6]. Moreover, $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is metrizable provided that $I$ is countable and the language $\mathcal{L}$ is separable (Definition 2.23); in particular, this holds for $\mathcal{L}_{\operatorname{tr}}$ [39, Observations 3.13 and 3.14]. See [32, §7] for further background. [^2] Technically, to make $\mathcal{L}_{\operatorname{tr}}$ separable by Definition 2.23, we need to have only countably many domains, and thus we should restrict the values of $r$ in the domains $D_{r}$ to a countable set such as $(0,\infty)\cap\mathbb{Q}$. However, the number of domains is irrelevant to metrizability of $\mathrm{S}_{\mathbf{D}}(\mathrm{T})$ for a fixed tuple of domains anyway. In the setting with domains of quantification, it is also convenient to have a topology on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$. We say that $\mathcal{O}\subseteq\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ is open if $\mathcal{O}\cap\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is open for every $\mathbf{D}\in\prod_{j\in\mathbb{N}}\mathcal{D}_{S_{j}}$; this defines a topology on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$, which we will also call the logic topology. One can then show that the inclusion map $\mathbb{S}_{\mathbf{D}}(\mathrm{T})\to\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ is a topological embedding [39, Observation 3.6]. Every formula defines a continuous function on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ for each theory $\mathrm{T}$, but the converse is not true. The objects that correspond to continuous functions on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ are a certain completion of the set of formulas, called definable predicates; see e.g. [32, §5.2]. Our approach to the definition will be semantic rather than syntactic, defining these objects immediately in terms of their interpretations.
@@
In this work, we will use definable predicates relative to $\mathrm{T}_{\operatorname{tr}}$, to $\mathrm{T}_{\mathrm{II}_{1}}$, and to an arbitrary complete $\mathrm{T}\models\mathrm{T}_{\operatorname{tr}}$. In [6, Theorem 9.9], the authors show in the setting with a single sort and a single domain that definable predicates relative to $\mathrm{T}$ correspond to continuous functions on the space of types $\mathrm{T}$, and this result was extended to the setting with multiple sorts and domains of quantification [39, Proposition 3.9]. Here we only need the following fact, for which we give a self-contained proof (the special case of $\mathrm{T}_{\operatorname{tr}}$ was done in [39, Remark 3.33]).
@@
Definable predicates relative to $\mathrm{T}$ have many of the same properties as formulas. Each definable predicate is uniformly continuous, where the same modulus of continuity serves as an upper bound in all models of $\mathrm{T}$; see [39, Observation 3.11]. The most important fact for this work is that definable predicates are closed under the same operations as formulas.
Fact 2.17 (See [6, Proposition 9.3] and [39, Lemma 3.12]).
@@ 2.6. Elementary extensions, saturation, and strong homogeneity
Now we are ready to formally define saturation. See also [22, §4.4] and [6, p. 35ff].
@@ 3.1. Definition and Properties
Note that $\Sigma$ is finitely approximately satisfiable in $\mathcal{M}$. Indeed, let $\delta>0$ and consider a finite collection of formulas $\phi_{1}$, …, $\phi_{n}\in\mathcal{E}$. Let $\phi=\max(|\phi_{1}|,\dots,|\phi_{n}|)$. By assumption, there exist infinitely many $y_{j}$’s satisfying $\phi^{\mathcal{M}}(y_{j})<\delta$ and $d(y_{j},y_{k})\geq\epsilon$ for $j\neq k$. This is enough to be able to approximately satisfy any finite collection of formulas from $\Sigma$. Since $|\Sigma|\leq\kappa^{\prime}<\kappa$ and $\mathcal{M}$ is $\kappa$-saturated, $\Sigma$ is satisfiable in $\mathcal{M}$. This means there exist $\kappa^{\prime}$ many $\epsilon$-separated elements $x_{j}\in D^{\mathcal{M}}$ such that $\phi^{\mathcal{M}}(x_{j})=0$ for all $\phi\in\mathcal{E}$, and hence $\operatorname{tp}^{\mathcal{M}}(x_{j}/A)=\operatorname{tp}^{\mathcal{M}}(x/A)$. Thus, the space of realizations of this type has density character greater than or equal to $\kappa^{\prime}$. Since $\kappa^{\prime}<\kappa$ was arbitrary, the its density character is at least $\kappa$. (4) $\implies$ (1). For each $k\in\mathbb{N}$, fix some formula $\phi_{k}$ and some $\delta_{k}>0$ such that $\phi_{k}^{\mathcal{M}}(x)=0$ and the set $\{y:\phi_{k}^{\mathcal{M}}(y)<\delta_{k}\}$ can be covered by finitely many $\epsilon$-balls. Let $C=\{y:\phi_{k}^{\mathcal{M}}(x)=0\text{ for all }k\in\mathbb{N}\}$. Then $C$ is a countable intersection of zerosets of definable predicates and hence is the zeroset of a definable predicate by [6, Proposition 9.14]. Note $C$ is closed (hence complete) by continuity of formulas. Moreover, by our choice of $\phi_{k}$, $C$ is totally bounded. Thus, $C$ is compact. By [6, Proposition 10.6], a compact zeroset is definable. Thus, $C$ is a compact definable set containing $x$, and so $x\in\operatorname{acl}^{\mathcal{M}}(A)$. (1) $\implies$ (5). If $x\in\operatorname{acl}^{\mathcal{M}}(A)$, then there exists a compact definable set $C$ such that $x\in C$. By [6, Proposition 10.4], $C$ (and hence $x$) must be contained in any elementary submodel of $\mathcal{M}$ that contains $A$. Note that this argument does not use saturation of $\mathcal{M}$. (5) $\implies$ (3). We proceed by contrapositive. Suppose that the space of realizations of $\operatorname{tp}^{\mathcal{M}}(x/A)$ in $\mathcal{M}$ has density character greater than or equal to $\kappa$. By the downward Löwenheim-Skolem theorem (Theorem 2.24), there exists an elementary submodel $\mathcal{P}$ of $\mathcal{M}$ that contains $A$ and has density character equal to that of $A$. Because the set of realizations of $\operatorname{tp}^{\mathcal{M}}(x/A)$ has strictly greater density character, there must be some realization $y$ of $\operatorname{tp}^{\mathcal{M}}(x/A)$ that is not contained in $\mathcal{P}$. Let $\mathbf{z}$ be a tuple on an index set $I$ of cardinality $\chi(A)$ so that $\mathbf{z}$ contains a dense subset of $A\cap D$ for each domain $D$. Then $(x,\mathbf{z})$ and $(y,\mathbf{z})$ have the same type because every $\mathcal{L}$-formula in $(x,\mathbf{z})$ is an $\mathcal{L}$-formula in $x$. Let $\mathbf{w}$ be a tuple with index set $I^{\prime}$ of cardinality less than $\kappa$ such that the elements of $\mathbf{w}$ in a domain $D$ are dense in $\mathcal{P}\cap D$; let $w_{i}$ be from the sort $S_{i}$ and domain $D_{i}$ for $i\in I^{\prime}$. We claim that there exists a tuple $\tilde{\mathbf{w}}$ on $I^{\prime}$ such that $\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\operatorname{tp}^{\mathcal{M}}(y,\mathbf{z},\mathbf{w})$. To see this, consider the set $\Phi$ of formulas $\phi$ with variables indexed by $\{0\}\sqcup I\sqcup I^{\prime}$ given by
@@
Let $B=(x,\mathbf{z})$ viewed as a set, and let $\Phi^{\prime}$ be the set of $L_{B}$ formulas on variables indexed by $I^{\prime}$, obtained from $\phi\in\Phi$ be plugging in $(x,\mathbf{z})$ for the first two arguments and viewing the formula as a function of the last argument. We claim that $\Phi^{\prime}$ is finitely approximately satisfiable in $\mathcal{M}$ as an $\mathcal{L}_{B}$ structure. To see this, consider a finite list of formulas $\phi_{1}^{\prime}$, …, $\phi_{n}^{\prime}$ from $\Phi^{\prime}$ corresponding to $\phi_{1}$, …, $\phi_{n}$ from $\Phi$. Note that $\phi_{1}$, …, $\phi_{n}$ only depend on a finite subset of the variables (by definition of formulas). Hence, it makes sense to define a formula in variables $(x^{\prime},\mathbf{z}^{\prime},\mathbf{w}^{\prime})$ by
@@
This demonstrates the finite approximate satisfiability of $\Phi^{\prime}$. Then since $\mathcal{M}$ is $\kappa$-saturated, $\Phi^{\prime}$ is satisfiable, so there exists $\tilde{\mathbf{w}}\in\prod_{i\in I^{\prime}}D_{i}^{\mathcal{M}}$ such that for $\phi^{\prime}\in\Phi$, we have $(\phi^{\prime})^{\mathcal{M}}(\tilde{\mathbf{w}})=0$. This means that for $\phi\in\Phi$, we have $\phi^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=0$. Note that for every formula $\phi$, there is some constant $c$ such that $\phi-c\in\Phi$. Therefore, we obtain that $\phi^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\phi^{\mathcal{M}}(y,\mathbf{z},\mathbf{w})$ for all formulas $\phi$, that is, $\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\mathbf{w})$. Let $\tilde{\mathcal{P}}$ be the substructure generated by $\tilde{\mathbf{w}}$, that is, the intersection of all substructures containing $\tilde{\mathbf{w}}$. We claim that for each sort $S$ and domain $D\in\mathcal{D}_{S}$, the set $\tilde{\mathcal{P}}\cap D^{\mathcal{M}}$ is the closure of the set $\Omega_{D}$ of elements of $\tilde{\mathbf{w}}$ that are in $D^{\mathcal{M}}$. It suffices to show that for each function symbol $f$ with inputs from sorts $S_{1}$, …, $S_{n}$ and output from sort $S$, if $f$ maps domains $D_{1}\times\dots\times D_{n}$ into domain $D$, then $f^{\mathcal{M}}$ maps $\overline{\Omega}_{D_{1}}\times\dots\times\overline{\Omega}_{D_{n}}$ into $\overline{\Omega}_{D}$. Furthermore, by continuity of $f^{\mathcal{M}}$ it suffices to show $f^{\mathcal{M}}$ maps $\Omega_{D_{1}}\times\dots\times\Omega_{D_{n}}$ into $\overline{\Omega}_{D}$. Let $\tilde{w}_{i_{1}}\in\Omega_{D_{1}}$, …, $\tilde{w}_{i_{n}}\in\Omega_{D_{n}}$ and let $\epsilon>0$. Note that $f^{\mathcal{M}}(w_{i_{1}},\dots,w_{i_{n}})$ is in $\mathcal{P}\cap D^{\mathcal{M}}$, so by our choice of $\mathbf{w}$, there exists some $i\in I$ such that
$d^{\mathcal{M}}(f^{\mathcal{M}}(\tilde{w}_{i_{1}},\dots,\tilde{w}_{i_{n}}),\tilde{w}_{i})=d^{\mathcal{M}}(f(w_{i_{1}},\dots,w_{i_{n}}),w_{i})<\epsilon,$
where the left equality follows because $w$ and $\tilde{w}$ have the same type. Since $\epsilon$ was arbitrary, $f^{\mathcal{M}}(\tilde{w}_{i_{1}},\dots,\tilde{w}_{i_{n}}))$ is in the closure of $\Omega_{D}$ as desired. Next, we claim $\tilde{\mathcal{P}}$ contains $A$. This is because for each domain $D$, each element of $\mathbf{z}$ in the domain $D$ is a limit of elements from $\mathbf{w}$ inside $D$, hence the same is true for $\tilde{\mathbf{w}}$ since $d^{\mathcal{M}}(z_{i},w_{j})=d^{\mathcal{M}}(z_{i},\tilde{w}_{j})$ by virtue of the types being the same. Similar reasoning shows that $\tilde{\mathcal{P}}$ does not contain $x$. Indeed, for each domain $D$, $d(x,D\cap\tilde{\mathcal{P}})$ is the infimum of distances between $x$ and $\tilde{w}_{i}$ for $i\in I^{\prime}$ such that $\tilde{w}_{i}\in D$. But $d(x,\tilde{w}_{i})=d(y,w_{i})$. Thus, $d(x,D\cap\tilde{\mathcal{P}}\cap D^{\mathcal{M}})=d(y,\mathcal{P}\cap D^{\mathcal{M}})>0$. Finally, we verify that $\tilde{\mathcal{P}}\preceq\mathcal{M}$ using the Tarski-Vaught test. Let $\phi$ be a formula in variables from sorts $S_{1}$, …, $S_{n+1}$, and let $D$ be a domain for $S_{n+1}$. Let $\tilde{w}_{i_{1}}$, …, $\tilde{w}_{i_{n}}$ be elements of $\tilde{w}$ from $S_{1}$, …, $S_{n}$ respectively. Since $\Omega_{D}$ is dense in $\tilde{\mathcal{P}}\cap D$, we have that
@@
Since the $\Omega_{D_{j}}$ is dense in $D_{j}^{\tilde{\mathcal{P}}}$ and both sides are continuous functions of the the first $n$ inputs, we obtain that for $\xi_{1}\in D_{1}^{\tilde{\mathcal{P}}}$, …, $\xi_{n}\in D_{n}^{\tilde{\mathcal{P}}}$,

### 09 Mar 2024 v3

@@ 2.4. Theories, models, and axioms
The class $\mathcal{C}$ is said to be axiomatizable if every $\mathcal{L}$-structure that models $\operatorname{Th}(\mathcal{C})$ is actually in $\mathcal{C}$. Farah, Hart, and Sherman [23, §3.2] showed that the class of $\mathcal{L}_{\operatorname{tr}}$-structures that represent actual tracial $\mathrm{W}^{*}$-algebras is axiomatizable. The theory of tracial $\mathrm{W}^{*}$-algebras will be denoted $\mathrm{T}_{\operatorname{tr}}$. Similarly, the class of $\mathrm{II}_{1}$-factors (infinite-dimensional tracial von Neumann algebras with trivial center) is axiomatizable and we denote its theory by $\mathrm{T}_{\mathrm{II}_{1}}$ [23].
@@ 2.5. Types, definable predicates, and definable functions
The space of types $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is equipped with a weak-*** topology as follows. If $\mathbf{S}$ is an $I$-tuple of $\mathcal{L}$-sorts, the set $\mathcal{F}_{\mathbf{S}}$ of formulas defines a real vector space. For each $\mathcal{L}$-structure $\mathcal{M}$ and $\mathbf{X}\in\prod_{i\in I}S_{j}^{\mathcal{M}}$, the type $\operatorname{tp}^{\mathcal{M}}(\mathbf{X})$ is a linear map $\mathcal{F}_{\mathbf{S}}\to\mathbb{R}$. Thus, for each $\mathcal{L}$-theory $\mathrm{T}$ and $\mathbf{D}\in\prod_{i\in I}\mathcal{D}_{S_{i}}$, the space $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is a subset of the real dual $\mathcal{F}_{\mathbf{S}}^{\dagger}$. We equip $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ with the weak-$\star$ topology (also known as the logic topology), which makes it into a compact Hausdorff space [6, Proposition 8.6]. Moreover, $\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is metrizable provided that $I$ is countable and the language $\mathcal{L}$ is separable (Definition 2.24); in particular, this holds for $\mathcal{L}_{\operatorname{tr}}$ [40, Observations 3.13 and 3.14]. See [33, §7] for further background. [^2] Technically, to make $\mathcal{L}_{\operatorname{tr}}$ separable by Definition 2.24, we need to have only countably many domains, and thus we should restrict the values of $r$ in the domains $D_{r}$ to a countable set such as $(0,\infty)\cap\mathbb{Q}$. However, the number of domains is irrelevant to metrizability of $\mathrm{S}_{\mathbf{D}}(\mathrm{T})$ for a fixed tuple of domains anyway. In the setting with domains of quantification, it is also convenient to have a topology on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$. We say that $\mathcal{O}\subseteq\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ is open if $\mathcal{O}\cap\mathbb{S}_{\mathbf{D}}(\mathrm{T})$ is open for every $\mathbf{D}\in\prod_{j\in\mathbb{N}}\mathcal{D}_{S_{j}}$; this defines a topology on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$, which we will also call the logic topology. One can then show that the inclusion map $\mathbb{S}_{\mathbf{D}}(\mathrm{T})\to\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ is a topological embedding [40, Observation 3.6]. Every formula defines a continuous function on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ for each theory $\mathrm{T}$, but the converse is not true. The objects that correspond to continuous functions on $\mathbb{S}_{\mathbf{S}}(\mathrm{T})$ are a certain completion of the set of formulas, called definable predicates; see e.g. [33, §5.2]. Our approach to the definition will be semantic rather than syntactic, defining these objects immediately in terms of their interpretations.
@@
In this work, we will use definable predicates relative to $\mathrm{T}_{\operatorname{tr}}$, to $\mathrm{T}_{\mathrm{II}_{1}}$, and to an arbitrary complete $\mathrm{T}\models\mathrm{T}_{\operatorname{tr}}$. In [6, Theorem 9.9], the authors show in the setting with a single sort and a single domain that definable predicates relative to $\mathrm{T}$ correspond to continuous functions on the space of types $\mathrm{T}$, and this result was extended to the setting with multiple sorts and domains of quantification [40, Proposition 3.9]. Here we only need the following fact, for which we give a self-contained proof (the special case of $\mathrm{T}_{\operatorname{tr}}$ was done in [40, Remark 3.33]).
@@
Definable predicates relative to $\mathrm{T}$ have many of the same properties as formulas. Each definable predicate is uniformly continuous, where the same modulus of continuity serves as an upper bound in all models of $\mathrm{T}$; see [40, Observation 3.11]. The most important fact for this work is that definable predicates are closed under the same operations as formulas.
Fact 2.17 (See [6, Proposition 9.3] and [40, Lemma 3.12]).
@@ 2.6. Elementary extensions, saturation, and strong homogeneity
Definition 2.20.
A substructure $\mathcal{M}\subseteq\mathcal{N}$ is an $\mathcal{L}$-structure where $S^{\mathcal{M}}\subseteq S^{\mathcal{N}}$ for each sort $S$, such that for each domain $D$ in a sort $S$, we have $D^{\mathcal{M}}=D^{\mathcal{M}}\cap S^{\mathcal{N}}$, and for each function symbol $f:S_{1}\times\dots\times S_{n}\to S$, we have $f^{\mathcal{M}}=f^{\mathcal{N}}|_{S_{1}^{\mathcal{M}}\times\dots\times S_{n}^{\mathcal{M}}}$.
In the case of von Neumann algebras, the reader may verify that the definition of substructure here coincides with the notion of von Neumann subalgebra, by using the Kaplansky density theorem to show that $D^{\mathcal{M}}=D^{\mathcal{N}}\cap S^{\mathcal{M}}$. In fact, this usage of the Kaplansky density theorem is already incorporated into the proof that von Neumann algebras can be axiomatized as $\mathcal{L}_{\operatorname{tr}}$-structures with the domains corresponding to operator norm balls [23, Proposition 3.3].
@@
Now we are ready to formally define saturation. See also [23, §4.4] and [6, p. 35ff].
@@ 3.1. Definition and Properties
Note that $\Sigma$ is finitely approximately satisfiable in $\mathcal{M}$. Indeed, let $\delta>0$ and consider a finite collection of formulas $\phi_{1}$, …, $\phi_{n}\in\mathcal{E}$. Let $\phi=\max(|\phi_{1}|,\dots,|\phi_{n}|)$. By assumption, there exist infinitely many $y_{j}$’s satisfying $\phi^{\mathcal{M}}(y_{j})<\delta$ and $d(y_{j},y_{k})\geq\epsilon$ for $j\neq k$. This is enough to be able to approximately satisfy any finite collection of formulas from $\Sigma$. Since $|\Sigma|\leq\kappa^{\prime}<\kappa$ and $\mathcal{M}$ is $\kappa$-saturated, $\Sigma$ is satisfiable in $\mathcal{M}$. This means there exist $\kappa^{\prime}$ many $\epsilon$-separated elements $x_{j}\in D^{\mathcal{M}}$ such that $\phi^{\mathcal{M}}(x_{j})=0$ for all $\phi\in\mathcal{E}$, and hence $\operatorname{tp}^{\mathcal{M}}(x_{j}/A)=\operatorname{tp}^{\mathcal{M}}(x/A)$. Thus, the space of realizations of this type has density character greater than or equal to $\kappa^{\prime}$. Since $\kappa^{\prime}<\kappa$ was arbitrary, its density character is at least $\kappa$. (4) $\implies$ (1). For each $k\in\mathbb{N}$, fix some formula $\phi_{k}$ and some $\delta_{k}>0$ such that $\phi_{k}^{\mathcal{M}}(x)=0$ and the set $\{y:\phi_{k}^{\mathcal{M}}(y)<\delta_{k}\}$ can be covered by finitely many $\epsilon$-balls. Let $C=\{y:\phi_{k}^{\mathcal{M}}(x)=0\text{ for all }k\in\mathbb{N}\}$. Then $C$ is a countable intersection of zerosets of definable predicates and hence is the zeroset of a definable predicate by [6, Proposition 9.14]. Note $C$ is closed (hence complete) by continuity of formulas. Moreover, by our choice of $\phi_{k}$, $C$ is totally bounded. Thus, $C$ is compact. By [6, Proposition 10.6], a compact zeroset is definable. Thus, $C$ is a compact definable set containing $x$, and so $x\in\operatorname{acl}^{\mathcal{M}}(A)$. (1) $\implies$ (5). If $x\in\operatorname{acl}^{\mathcal{M}}(A)$, then there exists a compact definable set $C$ such that $x\in C$. By [6, Proposition 10.4], $C$ (and hence $x$) must be contained in any elementary submodel of $\mathcal{M}$ that contains $A$. Note that this argument does not use saturation of $\mathcal{M}$. (5) $\implies$ (3). We proceed by contrapositive. Suppose that the space of realizations of $\operatorname{tp}^{\mathcal{M}}(x/A)$ in $\mathcal{M}$ has density character greater than or equal to $\kappa$. By the downward Löwenheim-Skolem theorem (Theorem 2.25), there exists an elementary submodel $\mathcal{P}$ of $\mathcal{M}$ that contains $A$ and has density character equal to that of $A$. Because the set of realizations of $\operatorname{tp}^{\mathcal{M}}(x/A)$ has strictly greater density character, there must be some realization $y$ of $\operatorname{tp}^{\mathcal{M}}(x/A)$ that is not contained in $\mathcal{P}$. Let $\mathbf{z}$ be a tuple of elements of $A$ on an index set $I$ of cardinality $\chi(A)$ so that $\mathbf{z}$ contains a dense subset of $A\cap D$ for each domain $D$. Then $(x,\mathbf{z})$ and $(y,\mathbf{z})$ have the same type because every $\mathcal{L}$-formula in $(x,\mathbf{z})$ is an $\mathcal{L}$-formula in $x$. Let $\mathbf{w}$ be a tuple of elements of $\mathcal{P}$ with index set $I^{\prime}$ of cardinality less than $\kappa$ such that the elements of $\mathbf{w}$ in each domain $D_{0}$ are dense in $\mathcal{P}\cap D^{\mathcal{M}}$; let $w_{i}$ be from the sort $S_{i}$ and domain $D_{i}$ for $i\in I^{\prime}$. We claim that there exists a tuple $\tilde{\mathbf{w}}$ indexed by $I^{\prime}$, of elements of $\mathcal{M}$ from appropriate sorts and domains, such that $\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\operatorname{tp}^{\mathcal{M}}(y,\mathbf{z},\mathbf{w})$. To see this, consider the set $\Phi$ of formulas $\phi$ with variables indexed by $\{0\}\sqcup I\sqcup I^{\prime}$ given by
@@
Let $B$ be the set of values taken by the tuple $(x,\mathbf{z})$, and let $\Phi^{\prime}$ be the set of $L_{B}$ formulas on variables indexed by $I^{\prime}$, obtained from $\phi\in\Phi$ be plugging in $(x,\mathbf{z})$ for the first two arguments and viewing the formula as a function of the last argument. We claim that $\Phi^{\prime}$ is finitely approximately satisfiable in $\mathcal{M}$ as an $\mathcal{L}_{B}$ structure. To see this, consider a finite list of formulas $\phi_{1}^{\prime}$, …, $\phi_{n}^{\prime}$ from $\Phi^{\prime}$ corresponding to $\phi_{1}$, …, $\phi_{n}$ from $\Phi$. Note that $\phi_{1}$, …, $\phi_{n}$ only depend on a finite subset of the variables (by definition of formulas). Hence, it makes sense to define a formula in variables $(x^{\prime},\mathbf{z}^{\prime},\mathbf{w}^{\prime})$ by
@@
This demonstrates the finite approximate satisfiability of $\Phi^{\prime}$. Then since $\mathcal{M}$ is $\kappa$-saturated, $\Phi^{\prime}$ is satisfiable, so there exists $\tilde{\mathbf{w}}\in\prod_{i\in I^{\prime}}D_{i}^{\mathcal{M}}$ such that for $\phi^{\prime}\in\Phi$, we have $(\phi^{\prime})^{\mathcal{M}}(\tilde{\mathbf{w}})=0$. This means that for $\phi\in\Phi$, we have $\phi^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=0$. Note that for every formula $\phi$, there is some constant $c$ such that $\phi-c\in\Phi$. Therefore, we obtain that $\phi^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\phi^{\mathcal{M}}(y,\mathbf{z},\mathbf{w})$ for all formulas $\phi$, that is, $\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\tilde{\mathbf{w}})=\operatorname{tp}^{\mathcal{M}}(x,\mathbf{z},\mathbf{w})$. Our goal in the rest of the proof is to show that the substructure $\tilde{\mathcal{P}}$ generated by $\tilde{\mathbf{w}}$ is an elementary substructure that contains $\mathbf{y}$ but not $x$. First, we give a concrete description of this substructure. For each sort $S_{0}$ and domain $D_{0}$, for each $w_{i}$ and $w_{j}$ in $D_{0}^{\mathcal{P}}$, we have that $d_{S_{0}}^{\mathcal{M}}(w_{i},w_{j})=d_{S_{0}}^{\mathcal{M}}(\tilde{w}_{i},\tilde{w}_{j})$, and in particular, if $w_{i}=w_{j}$, then $\tilde{w}_{i}=\tilde{w}_{j}$. We chose $\mathbf{w}$ such that the elements of $\mathbf{w}$ in a domain $D_{0}^{\mathcal{M}}$ are dense in $D_{0}^{\mathcal{P}}$. Since an isometry defined on a dense subset of a metric space extends to the entire metric space if the codomain is complete, there is an isometry $\alpha_{D_{0}}:D_{0}^{\mathcal{P}}\to D_{0}^{\mathcal{M}}$ such that $\alpha_{D_{0}}(w_{i})=\tilde{w}_{i}$ whenever $w_{i}\in D_{0}$. Moreover, for two domains $D_{0}$ and $D_{1}$ in the sort $S$, we have that $\alpha_{D_{0}}$ and $\alpha_{D_{1}}$ agree on $D_{0}^{\mathcal{P}}\cap D_{1}^{\mathcal{P}}$; to verify this, it suffices to show that $d_{S}^{\mathcal{M}}(\alpha_{D_{0}}(\xi),\alpha_{D_{1}}(\eta))=d_{S}^{\mathcal{M}}(\xi,\eta)$ holds for $\xi\in D_{0}^{\mathcal{P}}$ and $\eta\in D_{1}^{\mathcal{P}}$; this holds when $\xi$ and $\eta$ come from $\mathbf{w}$, by virtue of $\mathbf{w}$ and $\tilde{\mathbf{w}}$ having the same type, and then it extends to all of $D_{0}^{\mathcal{P}}$ and $D_{1}^{\mathcal{P}}$ by density. Define $\tilde{\mathcal{P}}$ by setting $D_{0}^{\tilde{\mathcal{P}}}:=\alpha_{D_{0}}(D_{0}^{\mathcal{P}})$ for each sort $S_{0}$ and domain $D_{0}$, and setting $S_{0}^{\tilde{\mathcal{P}}}=\bigcup_{D_{0}\in\mathcal{D}_{S_{0}}}D_{0}^{\tilde{\mathcal{P}}}$. We claim $\tilde{\mathcal{P}}$ is a substructure. First, $D^{\tilde{\mathcal{P}}}$ is complete because is the image of a complete metric space under an isometry. Second, we must show that $D_{0}^{\mathcal{M}}\cap S_{0}^{\tilde{\mathcal{P}}}=D_{0}^{\tilde{\mathcal{P}}}$ for each domain $D_{0}$; because $D_{0}^{\mathcal{M}}\cap S_{0}^{\mathcal{P}}=D_{0}^{\mathcal{P}}$, it suffices to check that for every domain $D_{1}$ for $\xi\in D_{1}^{\mathcal{P}}$, we have
$\inf_{\eta\in D_{0}^{\mathcal{M}}}d_{S}^{\mathcal{M}}(\alpha_{D_{1}}(\xi),\eta)=\inf_{\eta\in D_{0}^{\mathcal{M}}}d_{S}^{\mathcal{M}}(\xi,\eta);$
again this holds when $\xi$ is an element of $\mathbf{w}$ because $\mathbf{w}$ and $\tilde{\mathbf{w}}$ have the same type, and then it extends by continuity. Third, one shows that $\tilde{P}$ is closed under application of function symbols by verifying, in a similar manner to previous claims, that for each function symbol $f$ that maps domains $D_{1}\times\dots\times D_{n}$ into a domain $D_{0}$, we have
$d^{\mathcal{M}}(f(\alpha_{D_{1}}(\xi_{1}),\dots,\alpha_{D_{n}}(\xi_{n})),\alpha_{D_{0}}(\eta))=d^{\mathcal{M}}(f^{\mathcal{M}}(\xi_{1},\dots,\xi_{n}),\eta)\text{ for }\xi_{1}\in D_{1}^{\mathcal{P}},\dots,\xi_{n}\in D_{n}^{\mathcal{P}},\eta\in D_{0}^{\mathcal{P}},$
and hence $f^{\mathcal{M}}(\alpha_{D_{1}}(\xi_{1}),\dots,\alpha_{D_{n}}(\xi_{n}))=\alpha_{D_{0}}(f^{\mathcal{M}}(\xi_{1},\dots,\xi_{n}))$. Next, we verify that $\tilde{\mathcal{P}}\preceq\mathcal{M}$ using the Tarski-Vaught test. Let $\phi$ be a formula in variables from sorts $S_{1}$, …, $S_{n+1}$, and let $D$ be a domain for $S_{n+1}$. Let $\tilde{w}_{i_{1}}$, …, $\tilde{w}_{i_{n}}$ be elements of $\tilde{w}$ from $S_{1}$, …, $S_{n}$ respectively. Because the elements $w_{i}$ in a domain $D^{\mathcal{M}}$ are dense in $D^{\mathcal{P}}$, hence also the elements $\tilde{w}_{i}=\alpha_{D}(w_{i})$ in $D^{\mathcal{M}}$ are dense in $D^{\tilde{\mathcal{P}}}$, we have that
@@
Because the $\tilde{w}_{i}$’s that are in $D_{j}^{\tilde{\mathcal{P}}}$ form a dense subset of $D_{j}^{\tilde{\mathcal{P}}}$, and since both sides are continuous functions of the the first $n$ inputs, we obtain that for $\xi_{1}\in D_{1}^{\tilde{\mathcal{P}}}$, …, $\xi_{n}\in D_{n}^{\tilde{\mathcal{P}}}$,